# Fancy123: One Image to High-Quality 3D Mesh Generation via Plug-and-Play Deformation

### v1

@@ Abstract
Generating 3D meshes from a single image is an important but ill-posed task. Existing methods mainly adopt 2D multiview diffusion models to generate intermediate multiview images, and use the Large Reconstruction Model (LRM) to create the final meshes. However, the multiview images exhibit local inconsistencies, and the meshes often lack fidelity to the input image or look blurry. We propose Fancy123, featuring two enhancement modules and an unprojection operation to address the above three issues, respectively. The appearance enhancement module deforms the 2D multiview images to realign misaligned pixels for better multiview consistency. The fidelity enhancement module deforms the 3D mesh to match the input image. The unprojection of the input image and deformed multiview images onto LRM’s generated mesh ensures high clarity, discarding LRM’s predicted blurry-looking mesh colors. Extensive qualitative and quantitative experiments verify Fancy123’s SoTA performance with significant improvement. Also, the two enhancement modules are plug-and-play and work at inference time, allowing seamless integration into various existing single-image-to-3D methods.
Figure 1: Given a single image, our Fancy123 generates high-quality 3D meshes in terms of visual appeal, color clarity, and input fidelity.
@@ 1 Introduction
In summary, our contribution lies in proposing: (1) a new pipeline for the “123” task namely Fancy123, featuring two plug-and-play modules and highlighting the operation of unprojection to eliminate blurring effects; (2) an appearance enhancement module to correct multiview inconsistency via 2D image deformation; and (3) a fidelity enhancement module to improve the mesh’s similarity to the input image via 3D mesh deformation. Qualitative and quantitative experiments demonstrate that our Fancy123’s SoTA performance by significantly outperforming baseline methods. The effects of each module are verified in our ablation and backbone-replacement experiments. In particular, the two plug-and-play enhancement modules can be seamlessly integrated into various existing “123” methods.
@@ 2 Related work
Approaches for one-image-to-3D.  Benefiting from diffusion models [11] and large-scale 3D datasets [6], the task of reconstructing colored 3D mesh from a single image has made significant progress. The pioneer work DreamFusion [29] utilizes 2D diffusion models for 3D generation via score distillation sampling. To improve novel-view plausibility, multiview [32, 40] or view-conditioned [22] diffusion models are developed. With their generated multiview images as reference, 3D meshes can be generated via NeuS [41, 23], normal fusion [24], etc. To improve speed, researchers propose LRM [12], which is a feed-forward network trained with extensive 3D data for 3D generation, given single image or diffusion-generated multiview images as input. Besides, 3D diffusion models [19] or video diffusion models [47] are also used for 3D generation, yet not as popular as diffusion and LRM so far. While most existing methods yield blurry-looking meshes, Unique3D [45] achieves a clear appearance by unprojecting multiview images to the mesh, although its geometry reconstruction is slower and often yields artifacts compared to LRM. Approaches for multiview consistency.  To improve consistency among generated multiview images, existing methods mainly focus on network structures, training strategies, and sampling strategies of multiview diffusion models. MVDream [32] proposed a cross-view attention mechanism, and SyncDreamer [23] designed 3D-aware feature attention mechanism by unprojecting 2D features to a 3D volume. Zero123++ [31] argues for replacing Stable-Diffusion’s default scaled-linear schedule for sampling with linear schedule to emphasize global consistency over local details. Recently, ConsistNet [48] proposes a trainable plug-in block to enhance multiview consistency. Cycle3D [36] proposes to iteratively repeat 2D diffusion sampling and 3D feed-forward generation network. However, multiview inconsistency caused by pixel misalignment still exists.
@@ 3.1 Overview
(1) Initialization via multiview diffusion and LRM: We adopt a multiview diffusion model to generate high-quality multiview images $\{I_{k}\}$, based on which an LRM quickly obtains a corresponding initial mesh $\mathcal{M}_{0}$. We also follow [45, 24] to further refine the geometry by approximating diffusion-generated multiview normal maps.
@@
Our pipeline takes about 62 seconds on an NVIDIA A100 GPU when adopting InstantMesh as the backbone, with multiview diffusion 6 s, LRM 4 s, geometry refinement 27 s in step 1, 10 s in step 2, and 15 s in step 3, respectively. Note that geometry refinement can be skipped if LRM’s output mesh’s geometry is already satisfying. Effects of both modules are evaluated in Sec. 4.
@@ 3.2 Preliminaries
where $\mathbf{L}\in\mathbb{R}^{V\times V}$ is the Laplacian operator, $\boldsymbol{\nabla}$ is a stack of per-face gradient operators, and $\mathcal{A}\in\mathbb{R}^{3T\times 3T}$ is the mass matrix. For more details, please refer to APAP [49].
@@
Differentiable rendering. We denote the differentiable rendering process as $I^{R}=\mathcal{R}(\mathcal{M},\pi)$, where $\mathcal{M}$ is the input colroed mesh, $\pi$ is the camera parameters, and $I^{R}$ is the rendered image.
@@ 3.3 Initialization via multiview diffusion and LRM
Given an input image $I^{\text{in}}$, we generate multiview images $\{I_{k}\}$ by a multiview diffusion model, and then create a mesh by an LRM conditioned on $\{I_{k}\}$. Among existing open-source multi-view-input LMR, we adopt the latest InstantMesh [46] with its associated multiview diffusion model fine-tuned from [31], since it’s effective and fast. Other methods can also work. After that, we follow [45, 24] to further refine the mesh’s geometry, by optimizing mesh vertex coordiantes using multiview normal maps generated via  [31]. The output mesh of this step is denoted as $\mathcal{M}_{0}$.
@@ 3.4 Appearance enhancement
After several iterations, we can effectively alleviate the ghosting or blurry effect due to multiview inconsistency in the output mesh $\mathcal{M}_{c}$.
@@ 3.5 Fidelity enhancement
Camera pose estimation. The objective of 3D mesh deformation is to make the deformed mesh closely resemble the input image $I^{\text{in}}$ when observed from camera parameters $\pi$. Therefore, to reduce the difficulty of deformation, we should choose $\pi$ to ensure that the undeformed mesh $\mathcal{M}_{c}$ already has good correspondence with $I^{\text{in}}$. Some single-image-to-3D methods [45] use elevation and azimuth angles relative to $I^{\text{in}}$, so simply setting these two angles to zero can align $\mathcal{M}_{c}$ with $I^{\text{in}}$. However, our adopted IntantMesh [46] uses relative azimuth and absolute elevation angles. When $I^{\text{in}}$ has a big absolute elevation angle, such default camera setting may be unsuitable, necessitating camera pose estimation to find an appropriate $\pi$. Specifically, we perform a coarse-to-fine grid search over all possible elevation angles and select the $\pi$ that minimizes the LPIPS [50] score: $\text{LPIPS}(I^{\text{in}},\mathcal{R}(\mathcal{M}_{c},\pi))$, where LPIPS is a metric indicating the difference between two images. We also add a small optimization loop to further optimize $\pi$. See our supplementary file for more details.
@@
c—ccccc—cc Appearence Metrics   Geometry Metrics
@@
Unproject input image. After the 3D deformation, we can safely unproject $I^{\text{in}}$ to $\mathcal{M}_{d}$ to obtain the final output mesh $\mathcal{M}_{\text{output}}=\mathcal{P}(\mathcal{M}_{d},\{I^{\text{in}}\},\{\pi\})$.
@@ 4.1 Comparison with SoTA methods
As mentioned, some methods (InstantMesh and LGM) do not align with the input view by default, so we align them by a coarse-to-fine grid search, and adopt the elevation and azimuth with the lowest LPIPS score for rendering. See our supplementary for more details.
@@
However, we experimentally find that comparing GT and generated meshes using these metrics is not reliable, as it does not align well with human perception. For example, Fig. 5 (b-d) shows increasing appearance quality by human perception, but the PSNR scores are 12.76, 12.35, and 12.44, respectively, indicating (b) is the best and (c) is the worst. Other metrics also have the similar issue; see our supplementary file for more examples. We analyze that the inconsistency between quantitative metrics and visual judgments is mainly due to: (1) All current metrics solely assess fidelity (i.e., the similarity between GT and generated mesh), not plausibility. If two meshes both poorly match the input, the more plausible one deserves a higher score, but existing metrics treat them equally. (2) Due to the input’s lack of perspectives, the GT of this task should be diverse. For example, in Fig. 4, any color on the back of the red bag is reasonable as long as there are no obvious artifacts. Therefore, evaluating with a single definite GT is improper.
@@ 4.2 Ablation study
Figure 7: Improvement is significant when applying our various plug-and-play modules to other backbones.
@@
(g) Directly unprojecting input images to the mesh causes artifacts (lack of an ear, too wide of an eye, etc.).
@@
Backbone replacement. Besides InstanceMesh, our appearance and fidelity enhancement modules can be used in other backbones. In Fig. 7, significant improvement can be seen when applying (a) our fidelity enhancement to TripoSR [38], (b) our appearance enhancement to Unique3D [45], and (c) our complete pipeline to LGM [35].
@@ 5 Conclusion
We propose Fancy123, a novel SoTA framework for the single-image-to-3D-mesh task. Fancy123 features appearance and fidelity enhancement modules to enhance multiview consistency and fidelity by plug-and-play deformation, and an unprojection operation to ensure clear coloration. Though promising reconstruction quality has been verified in extensive experiments, our Fancy123 still has limitations. First, like other methods, Fancy123’s performance is constrained by the adopted multiview diffusion model. Second, our 3D deformation is only supervised with RGBA images, thus can sometimes lead to artifacts when the object’s different semantic parts share similar colors. In the future, we plan to investigate adopting semantic guidance to address this issue, and expand our method to adapt to large-scale scenes.
@@ Appendix B More Quantitative results
c—ccccc—cc Appearence Metrics   Geometry Metrics
@@ B.2 Inadequacies of existing quantitative metrics
Appendix C More qualitative results
@@
C.1 Comparison with SoTA
We provide more visual comparisons between our Fancy123 and baseline methods in Figs. 11 and 12. InstantMesh, CRM, TripoSR, and LGM often produce blurry-looking meshes. SF3D and TripoSR exhibit lower plausibility in novel views. Unique3D often yields significant artifacts as shown in the red boxes. Unlike them, our Fancy123 archives high clarity and plausibility for various challenging input images.
C.2 Ablation study
2D deformation. Figs. 13 and 14 present more visual results on the effect of our 2D-deformation-based appearance enhancement module.

### v2

@@ Abstract
Generating 3D meshes from a single image is an important but ill-posed task. Existing methods mainly adopt 2D multiview diffusion models to generate intermediate multiview images, and use the Large Reconstruction Model (LRM) to create the final meshes. However, the multiview images exhibit local inconsistencies, and the meshes often lack fidelity to the input image or look blurry. We propose Fancy123, featuring two enhancement modules and an unprojection operation to address the above three issues, respectively. The appearance enhancement module deforms the 2D multiview images to realign misaligned pixels for better multiview consistency. The fidelity enhancement module deforms the 3D mesh to match the input image. The unprojection of the input image and deformed multiview images onto LRM’s generated mesh ensures high clarity, discarding LRM’s predicted blurry-looking mesh colors. Extensive qualitative and quantitative experiments verify Fancy123’s SoTA performance with significant improvement. Also, the two enhancement modules are plug-and-play and work at inference time, allowing seamless integration into various existing single-image-to-3D methods. Project page: https://github.com/YuQiao0303/Fancy123.
Figure 1: Given a single image, our Fancy123 generates a high-quality 3D mesh in terms of visual appeal, color clarity, and input fidelity.
@@ 1 Introduction
In summary, our contribution lies in proposing: (1) a new pipeline namely Fancy123 for enhancing “123” generation quality, featuring two plug-and-play modules and highlighting the operation of unprojection to avoid blurriness; (2) an appearance enhancement module to correct multiview inconsistency via 2D image deformation; and (3) a fidelity enhancement module to improve the mesh’s similarity to the input image via 3D mesh deformation. Qualitative and quantitative experiments demonstrate that our Fancy123’s SoTA performance by significantly outperforming baseline methods. The effects of each module are verified in our ablation and backbone-replacement experiments. In particular, the two plug-and-play enhancement modules can be seamlessly integrated into various existing “123” methods.
@@ 2 Related work
Approaches for one-image-to-3D.  Benefiting from diffusion models [11] and large-scale 3D datasets [6], the task of reconstructing colored 3D mesh from a single image has made significant progress [54, 29]. The pioneer work DreamFusion [31] utilizes 2D diffusion models for 3D generation via score distillation sampling. To improve novel-view plausibility, multiview [34, 44] or view-conditioned [23] diffusion models are developed. With their generated multiview images as reference, 3D meshes can be generated via NeuS [45, 24], normal fusion [25], etc. To improve speed, researchers propose LRM [12], which is a feed-forward network trained with extensive 3D data for 3D generation, given single image or diffusion-generated multiview images as input. Besides, 3D diffusion models [19] or video diffusion models [51] are also used for 3D generation, yet not as popular as diffusion and LRM so far. While most existing methods yield blurry-looking meshes, Unique3D [49] achieves a clear appearance by unprojecting multiview images to the mesh, although its geometry reconstruction is slower and often yields artifacts compared to LRM.
Approaches for multiview consistency.  To improve consistency among generated multiview images, existing methods mainly focus on network structures, training strategies, and sampling strategies of multiview diffusion models. MVDream [34] proposed a cross-view attention mechanism, and SyncDreamer [24] designed a 3D-aware feature attention mechanism by unprojecting 2D features to a 3D volume. Zero123++ [33] argues for replacing Stable-Diffusion’s default scaled-linear schedule for sampling with a linear schedule to emphasize global consistency over local details. Recently, ConsistNet [52] proposes a trainable plug-in block to enhance multiview consistency. Cycle3D [40] proposes to iteratively repeat 2D diffusion sampling and 3D feed-forward generation network. However, multiview inconsistency caused by pixel misalignment still exists.
@@ 3.1 Overview
(1) Initialization via multiview diffusion and LRM: We adopt an multiview diffusion model to generate high-quality multiview images $\{I_{k}\}$, based on which an LRM quickly obtains a corresponding initial mesh $\mathcal{M}_{0}$. We also follow [49, 25] to further refine the geometry by approximating diffusion-generated multiview normal maps.
@@
Our pipeline takes about 62 seconds on an NVIDIA A100 GPU when adopting InstantMesh as the backbone, with multiview diffusion 6 s, LRM 4 s, geometry refinement 27 s in step 1, 10 s in step 2, and 15 s in step 3, respectively. The peak GPU memory usage is about 21M.
@@ 3.2 Preliminaries
[…]
@@
Differentiable rendering. We denote the differentiable rendering process as $I^{R}=\mathcal{R}(\mathcal{M},\pi)$, where $\mathcal{M}$ is the input colored mesh, $\pi$ is the camera parameters, and $I^{R}$ is the rendered image.
@@ 3.3 Initialization via multiview diffusion and LRM
Given an input image $I^{\text{in}}$, we generate multiview images $\{I_{k}\}$ by a multiview diffusion model, and then create a mesh by an LRM conditioned on $\{I_{k}\}$. Among existing open-source multi-view-input LRM, we adopt the latest InstantMesh [50] with its associated multiview diffusion model fine-tuned from [33], since it’s effective and fast. Other methods can also work. After that, we follow [49, 25] to further refine the mesh’s geometry, by optimizing mesh vertex coordiantes using multiview normal maps generated via  [33]. The output mesh of this step is denoted as $\mathcal{M}_{0}$.
@@ 3.4 Appearance enhancement
After 100 (experimentally determined) iterations, we can effectively alleviate the ghosting or blurry effect due to multiview inconsistency in the output mesh $\mathcal{M}_{c}$.
@@ 3.5 Fidelity enhancement
Camera pose estimation. The objective of 3D mesh deformation is to make the deformed mesh closely resemble the input image $I^{\text{in}}$ when observed from camera parameters $\pi$. Therefore, to reduce the difficulty of deformation, we should choose $\pi$ to ensure that the undeformed mesh $\mathcal{M}_{c}$ already has good correspondence with $I^{\text{in}}$. Some single-image-to-3D methods [49] use elevation and azimuth angles relative to $I^{\text{in}}$, so simply setting these two angles to zero can align $\mathcal{M}_{c}$ with $I^{\text{in}}$. However, our adopted IntantMesh [50] uses relative azimuth and absolute elevation angles. When $I^{\text{in}}$ has a big absolute elevation angle, such default camera setting may be unsuitable, necessitating camera pose estimation to find an appropriate $\pi$. Specifically, we perform a coarse-to-fine parameter search over all possible elevation angles and select the $\pi$ that minimizes the LPIPS [55] score: $\text{LPIPS}(I^{\text{in}},\mathcal{R}(\mathcal{M}_{c},\pi))$, where LPIPS is a metric indicating the difference between two images. We also add a small optimization loop to further optimize $\pi$. See our supplementary file for more details.
@@
c—ccccc—$c$ Appearence Metrics   Geometry Metrics
@@
Unproject input image. After the 3D deformation (200 iterations), we can safely unproject $I^{\text{in}}$ to $\mathcal{M}_{d}$ to obtain the final output mesh $\mathcal{M}_{\text{output}}=\mathcal{P}(\mathcal{M}_{d},\{I^{\text{in}}\},\{\pi\})$.
@@ 4.1 Comparison with SoTA methods
As mentioned, some methods (InstantMesh and LGM) do not align with the input view by default, so we align them by coarse-to-fine parameter search, and adopt the camera parameters with the lowest LPIPS score for rendering. See our supplementary for more details.
@@
However, we experimentally find that comparing GT and generated meshes using these metrics is not reliable, as it does not align well with human perception. For example, Fig. 5 (b-d) shows increasing appearance quality by human perception, but the PSNR scores are 12.76, 12.35, and 12.44, respectively, indicating (b) is the best and (c) is the worst. Other metrics also have the similar issue; see our supplementary file for more examples. We analyze that the inconsistency between quantitative metrics and visual judgments is mainly due to: (1) All current metrics solely assess fidelity (i.e., the similarity between GT and generated mesh), not plausibility. If two meshes both poorly match the input, the more plausible one deserves a higher score, but existing metrics treat them equally. (2) Due to the input’s lack of perspectives, the GT of this task should be diverse. For example, in Fig. 4, any color on the back of the red bag is reasonable as long as there are no obvious artifacts. Therefore, evaluating with a single definite GT is inproper.
@@ 4.2 Ablation study
Figure 7: Improvement is significant when applying our various plug-and-play enhancement modules to other backbones.
@@
(g) directly unprojecting input image to the mesh causes artifacts (lack of an ear, too wide eye etc.)
@@
Backbone replacement. Besides InstanceMesh, our proposed appearance and fidelity enhancement modules can be flexibly used in other backbone methods. In Fig. 7, significant improvement can be seen when applying (a) our fidelity enhancement to TripoSR [42], (b) our appearance enhancement to Unique3D [49], and (c) our complete refinement pipeline to LGM [38].
@@ 5 Conclusion
We propose Fancy123, a novel SoTA framework for enhancing single-image-to-3D-mesh generation quality. It features appearance and fidelity enhancement modules to enhance multiview consistency and fidelity by plug-and-play deformation, and an unprojection operation to ensure clear coloration. Though promising reconstruction quality has been verified in extensive experiments, Fancy123 still has limitations. First, like other methods, its performance is constrained by the adopted multiview diffusion model. Second, our 3D deformation is only supervised with RGBA images, thus can sometimes cause artifacts when the object’s different semantic parts share similar colors. See our supplementary file for an example. In the future, we plan to investigate adopting semantic guidance to address this issue, and expand our method to adapt to large-scale scenes.
Acknowledgements
This work was supported by the China National Natural Science Foundation No. 62202182, Guangdong Basic and Applied Basic Research Foundation 2024A1515010224, and also by the National Key Research and Development Program of China under Grant 2023YFB4503400, the China National Natural Science Foundation No.62176101, 62276109, 62450064, and 62322205.
@@ Appendix B More Quantitative results
c—ccccc—$c$ Appearence Metrics   Geometry Metrics
@@ B.2 Inadequacies of existing quantitative metrics
Figure 10: (a)-(b) Applying our Fancy123’s enhancement modules to different initialization methods. (b) Fancy123 improves the results even for relatively low-quality initial meshes. (c) Comparison of Fancy123’s refinement performance against Other refinement techniques.
Appendix C Initialization method and enhancement module replacement experiment
To validate the universal applicability of the enhancement modules in Fancy123, we replace the mesh initialization method InstantMesh with other baseline methods, and apply Fancy123 enhancement modules as follows:
(1) TripoSR/SF3D + our enhancement: Since neither TripoSR nor SF3D generates multiview images as intermediate products, Fancy123’s appearance enhancement module can not be applied. Thus, only geometry and fidelity enhancements are applied.
(2) Unique3D + our enhancement: Since Unique3D already incorporates geometry enhancement identical to Fancy123, only appearance and fidelity enhancements are applied.
(3) LGM/CRM/InstantMesh + our enhancement: We apply out complete enhancement pipeline (geometry, appearance, and fidelity) to meshes generated by LGM, CRM, and InstantMesh. For LGM and CRM, we generate multiview normal maps by the normal estimation model provided by Unique3D. All these three methods use the same four viewpoints (front, back, left, right), ensuring good model compatibility. As shown in Fig. 10 (a) and (b), our enhancement modules improved the generation quality for all initialization methods. Even for low-quality initial meshes in  Fig. 10 (b) with geometric flaws (e.g., marked by red boxes ), Fancy123 enhances coherence, smoothness, and overall plausibility.
Additionally, we compare Fancy123 with two typical enhancement techniques for single-image-to-3D:
(1) Input-MSE Enhancement: This intuitive approach optimizes mesh parameters by minimizing the mean squared error (MSE) loss between input images and mesh renderings. It is widely adopted in image-conditioned 3D generation (e.g., Make-It-3D [37], DreamGaussian [39]), and it only enhances regions visible from the input viewpoint. However, it produces significant artifacts when geometric misalignment exists between the mesh and input image, as shown in  Fig. 10 (c) “Input-MSE refine’. In contrast, our Fancy123 achieves precise alignment through 3D deformation, avoiding such artifacts.
(2) SDS/PDM (Perturb-Denoise-MSE) Enhancement: SDS [31] is a foundational method for ”2D diffusion for 3D generation”. It’s also commonly used in refinement stages [21]. It works by manually adding noise to mesh renderings and optimizing mesh parameters via SDS loss. DreamGaussian [39] notes that SDS-based UV color optimization introduces unwanted artifacts and proposes to replace SDS loss with MSE after perturbation, which we term “PDM”. While both SDS and PDM enhance the whole meshes across all viewpoints, they suffer from texture blurring due to inconsistent 3D guidance signals inherent in diffusion denoising, as mentioned by DreamGaussian [39] authors. As shown in  Fig. 10 (c) “PDM refine”, PDM may yield an even blurrier texture than the initial mesh.
Quantitative results in Tab. 3 show that our Fancy123 enhances all baselines, while Input-MSE and PDM degrade most metrics (green/red text indicating performance improvement/decline). We assume the key reasons are as mentioned above: (1) Input-MSE fails to handle geometric misalignment between 3D meshes and input images. (2) PDM reduces texture clarity due to inconsistent denoising guidance.
FID ↓ LPIPS ↓ PSNR ↑ SSIM ↑ CLIP-Sim. ↑
LGM 86.24 0.395 13.254 0.612 0.710
+our refine 58.92 0.377 13.266 0.610 0.748
+input-MSE refine 69.44 0.407 12.745 0.575 0.696
+PDM refine 102.80 0.397 13.163 0.599 0.669
CRM 87.74 0.388 13.664 0.620 0.729
+our refine 57.65 0.368 13.756 0.626 0.784
+input-MSE refine 75.31 0.389 13.669 0.590 0.725
+PDM refine 102.18 0.391 13.570 0.594 0.686
TripoSR 68.52 0.367 13.876 0.629 0.739
+our refine 51.76 0.347 14.144 0.635 0.780
+input-MSE refine 66.96 0.369 13.880 0.619 0.726
+PDM refine 89.82 0.376 13.934 0.591 0.703
InstantMesh 46.16 0.349 13.735 0.634 0.800
+our refine 37.99 0.330 14.365 0.651 0.835
+input-MSE refine 40.82 0.354 13.848 0.619 0.803
+PDM refine 72.94 0.377 13.837 0.596 0.734
Unique3D 58.51 0.389 12.741 0.592 0.764
+our refine 53.69 0.379 13.094 0.597 0.774
+input-MSE refine 57.66 0.396 12.867 0.583 0.742
+PDM refine 97.35 0.414 12.918 0.558 0.662
SF3D 49.89 0.343 14.320 0.617 0.776
+our refine 50.66 0.332 15.091 0.648 0.790
+input-MSE refine 49.71 0.348 14.685 0.615 0.769
+PDM refine 79.77 0.359 14.220 0.621 0.728
Table 3: Results of different initialization methods without or with different refinement methods. Green and red indicate better and worse performance after refinement, respectively.
Appendix D Can the order of enhancement modules be swapped?
No. (1) The initial mesh without appearance enhancement (AE) may be too blurry for fidelity enhancement (FE), since 3D deformation in fidelity enhancement is guided by RGB loss. (2) AE unprojects the deformed multiview images onto the mesh, and FE unprojects the input image. If we first apply FE and then AE, the result of FE will be overwritten by AE. So we recommend following the order in our paper to apply AM then FM, so as to first address multiview inconsistency (across all views) by AE then improve the input view quality by FE.
Appendix E Failur case
@@ Appendix F More qualitative results
F.1 Comparison with SoTA
We provide more visual comparisons between our Fancy123 and baseline methods in Figs. 12 and 13. InstantMesh, CRM, TripoSR, and LGM often produce blurry-looking meshes. SF3D and TripoSR exhibit lower plausibility in novel views. Unique3D often yields significant artifacts as shown in the red boxes. Unlike them, our Fancy123 archives high clarity and plausibility for various challenging input images.
F.2 Ablation study
2D deformation. Figs. 14 and 15 present more visual results on the effect of our 2D-deformation-based appearance enhancement module.